OpenStudy (anonymous):

Find two consecutive integers whose sum is 113. Which of the following equations could be used to solve the problem?

OpenStudy (raden):

let that numbers are x, x+1 respectively. so we have an equation : x + x + 1 = 113 solve for x, first

OpenStudy (anonymous):

100 and 12? @RadEn

OpenStudy (raden):

100 and 12 ? consecutive numbers like 12 and 13, 15 and 16, 99 and 100, 100 and 101

OpenStudy (anonymous):

Ohh. 99 and 13?

OpenStudy (raden):

just solve this with of algebra x + x + 1 = 113, can you ?

OpenStudy (anonymous):

BTW, here are the choices I can choose from. 2x = 113 2x + 1 = 113 2x + 2 = 113 x2 + 1= 113

OpenStudy (raden):

it is the second one, combine the similar terms x + x + 1 = 113 2x + 1 = 113

OpenStudy (raden):

now solve for x

OpenStudy (anonymous):

Makes so much sense now! Is it 56?

OpenStudy (raden):

hehe, correct. x = 56 and the second number is x + 1 = 56 + 1 = 57 s0, they are 56 and 57 resepectively

OpenStudy (anonymous):

Makes soo much sense!! :D Thanks a bunch! You rock!

OpenStudy (raden):

welcome :)